# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v2

@@ 4.1 Datasets
The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024, ), permitted strictly for non-commercial editorial use (International Olympic Committee, ). These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data. This dataset allows us to assess how well our approach adapts to newly introduced visual styles that the base model has never encountered.
@@ 4.2 Style Transfer Results
Figure 3: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of $<$@eveismyname$>$”, with the embedding $<$@eveismyname$>$ trained using different values of $m$ and $\varepsilon$.
Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using different values of $m$ and $\varepsilon$.
@@
Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al., 2023) embeddings on Stable Diffusion v1.5 (Rombach et al., 2022) using DPAgg-TI. Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience. For regular TI, we utilize the default process to embed the private dataset without any additional noise. For the DPAgg-TI, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy.
@@
Figures 3 and  4 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$. We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparision between different configurations. As with common practice, we set $\delta=1/n$. Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$. The purpose of this parameter value is to demonstrate the image generated when $u^{*}$ contains zero information about the target dataset. Images generated without DP closely resemble the unique stylistic elements of the target dataset. In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure. In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in more noticeable noise and diminished stylistic fidelity.
@@
As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity. In particular, $y_{u^{*}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding). In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms). With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees. We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 3 and intermediate $\varepsilon$ values in Figure 4) which restores as $u^{*}$ drifts even further from its optimal value. We hypothesize that this is due to drifted $u^{*}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}$ to be disregarded from $y_{u^{*}}$ and the prompt fidelity to be restored. Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}$ passing through non-linear regions within embedding space. We leave further investigations into this observation for future work.
@@
Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$. This strong boost in robustness comes at a small price of base style capture fidelity. As observed in Figures 3 and 4, we can also treat subsampling as an introduction of noise. Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid. However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the biggest outlier in the dataset.
@@ 4.3.1 User Study
4.3.2 Kernel Inception Distance
@@
The Kernel Inception Distance (KID) (Bińkowski et al., 2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space. To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al., 2015), and their distributional differences are estimated. Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al., 2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al., 2024), making it more suitable for smaller datasets, as in our case.
@@
We report KID scores for different parameters in Tables 2 and 2, showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy. Further discussion of these results is provided in Appendix B.
Table 1: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
0.0444 0.0794 0.0422 0.0529 0.0690 0.1117 0.0654
32 0.0752 0.0845 0.0865 0.1167 0.0300 0.0649 0.0657
16 0.0351 0.0379 0.0430 0.0663 0.1309 0.0438 0.0658
8 0.0359 0.0366 0.0350 0.0366 0.0396 0.0530 0.0658
4 0.0245 0.0250 0.0249 0.0250 0.0258 0.0314 0.0653
ctrl 0.0318
Table 2: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
0.1146 0.1202 0.1368 0.1314 0.1389 0.1209 0.1274
32 0.1220 0.1036 0.1258 0.1377 0.1307 0.1245 0.1259
16 0.1311 0.1424 0.1170 0.1311 0.1381 0.1335 0.1278
8 0.1317 0.1307 0.1220 0.1117 0.1295 0.1313 0.1272
4 0.1141 0.1094 0.1137 0.1190 0.1194 0.1583 0.1259
ctrl 0.1388
4.4 Ablation Study
4.4.1 Textual Inversion with DP-SGD
Figure 5: Comparing our approach to applying DP-SGD to regular TI using prompt “a painting of Taylor Swift in the style of @eveismyname”.
@@ 4.4.1 Textual Inversion with DP-SGD
A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training. We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets. We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset. In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 5. We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images. Additional results can be found in Appendix D.
4.4.2 Differentially Private Adaptation Using Style Guidance
We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al., 2024). Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
We encode each target image $x^{(i)}$ as $u^{(i)}$ via a CLIP image encoder, then aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ into $u^{*}$ using (5) or (7), depending on whether subsampling is applied. The aggregated embedding $u^{*}$ is then incorporated into the reverse diffusion process as a style guide. Further implementation detailes are provided in Appendix C.
@@ 5 Conclusion
We presented a differentially private adaptation method for diffusion models using Textual Inversion for privacy-preserving style transfer. Experiments on private artwork and Paris 2024 pictograms showed TI preserves stylistic fidelity and outperforms Style Guidance. Our results demonstrate embedding-driven methods as efficient, scalable alternatives to DP-SGD, balancing style quality and privacy.
Impact Statement
@@
Acknowledgements
We sincerely thank Tatchamon Wongworakul (@eveismyname) for providing her artwork for use in this study. We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study. Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
@@ A.5 Results and Analysis
Table 3: Survey Results.
regular TI No Adaptation Unsure
@eveismyname 19 4 2
Paris 2024 16 6 3
DPAgg-TI (no DP, no subsampling) No Adaptation Unsure
@eveismyname 16 9 0
Paris 2024 15 4 6
regular TI DPAgg-TI (no DP, no subsamp.) Unsure
@eveismyname 12 13 0
Paris 2024 9 10 6
regular TI DPAgg-TI (no DP, subsamp. $m=8$) Unsure
@eveismyname 16 6 3
Paris 2024 7 13 5
@eveismyname 18 4 3
Paris 2024 10 8 7
@eveismyname 14 10 1
Paris 2024 3 16 6
DPAgg-TI (no DP, no subsampling) Style Guidance Unsure
@eveismyname 16 8 1
Paris 2024 20 2 3
DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Style Guidance Unsure
@eveismyname 16 8 1
Paris 2024 19 2 4
@eveismyname 8 5 12
Paris 2024 15 4 6
Appendix B Kernel Inception Distance Discussion
@@
For KID evaluations, we used prompts similar to those employed during TI training: “A painting/icon in the style of $S^{*}$”. Consistent with the training image captions, these prompts do not specify a subject.
@@
Appendix C Style Guidance
@@ C.1 Background: Denoising Diffusion Implicit Models
Denoising Diffusion Implicit Models (DDIM) sampling (Song et al., 2021a) uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$. The reverse diffusion process for DDIM sampling can be formalized as follows:

### v3

@@ 4.1 Datasets
The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024 2023), permitted strictly for non-commercial editorial use (International Olympic Committee 2025). These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data. This dataset allows us to assess how well our approach adapts to newly introduced visual styles that the base model has never encountered.
@@ 4.2 Style Transfer Results
Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al. 2023) embeddings on Stable Diffusion v1.5 (Rombach et al. 2022) using DPAgg-TI. Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience. For regular TI, we utilize the default process to embed the private dataset without any additional noise. For the DPAgg-TI, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy.
@@
Figures 4 and  5 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$. We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparison between different configurations. As with common practice, we set $\delta=1/n$. Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$. The purpose of this parameter value is to demonstrate the image generated when $u^{*}_{\text{DP}}$ contains zero information about the target dataset. Images generated without DP closely resemble the unique stylistic elements of the target dataset. In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure. In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in diminished stylistic fidelity.
@@
As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}_{\text{DP}}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity. In particular, $y_{u^{*}_{\text{DP}}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding). In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms). With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}_{\text{DP}}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees. We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 4 and intermediate $\varepsilon$ values in Figure 5) which restores as $u^{*}_{\text{DP}}$ drifts even further from its optimal value. We hypothesize that this is due to drifted $u^{*}_{\text{DP}}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}_{\text{DP}}$ to be disregarded from $y_{u^{*}_{\text{DP}}}$ and the prompt fidelity to be restored. Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}_{\text{DP}}$ passing through non-linear regions within embedding space. We leave further investigations into this observation for future work.
@@
Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$. This strong boost in robustness comes at a small price of base style capture fidelity. As observed in Figures 4 and 5, we can also treat subsampling as an introduction of noise. Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid. However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the biggest outlier in the dataset.
@@ 4.3 Quantitative Evaluation
Kernel Inception Distance
@@
The Kernel Inception Distance (KID) (Bińkowski et al. 2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space. To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al. 2015), and their distributional differences are estimated. Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al. 2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al. 2024), making it more suitable for smaller datasets, as in our case.
@@
We report KID scores for different parameters in Tables  2 and 3 (see Appendix B), showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy. Further discussion of these results is also provided in Appendix B.
4.4 Ablation Study: Textual Inversion with DP-SGD
@@
A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training. We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets. We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset. In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 6. We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images. Additional results can be found in Appendix F.
Figure 6: Comparing our approach to applying DP-SGD to regular TI using prompts “an icon of a dragon in the style of the Paris 2024 Olympic Pictograms” and “a painting of Taylor Swift in the style of @eveismyname” respectively. Note that our method aggregates individual TI embeddings for each training image, whereas the baseline trains a single TI embedding over the entire dataset.
5 Copyright Protection Implications
Our proposed mechanism can also be interpreted through the lens of copyright protection. This connection is grounded in the framework of Near Access-Freeness (NAF) (Vyas, Kakade, and Barak 2023), which evaluates whether a model’s outputs reveal undue influence from specific data points by comparing them to those from a safe model trained without access to the same data.
Since DPAgg-TI satisfies $\varepsilon$-DP, it also satisfies $\varepsilon$-NAF, which means the adapted model behaves similarly to one that never saw the private images, under the NAF criterion. However, we emphasize that this guarantee holds only within the NAF framework; it does not constitute a general claim about content similarity or legal compliance. Crucially, DPAgg-TI is designed to adapt to the style of private images, not their specific content. Prior work and legal precedent suggest that style imitation is generally considered fair use and does not constitute infringement (Vyas, Kakade, and Barak 2023). Thus, our mechanism aligns with the intended protections of NAF: it avoids memorization while still enabling meaningful personalization and stylistic adaptation. We defer the details of copyright protection to Appendix D.
@@ 6 Conclusion
We presented a differentially private adaptation method for diffusion models based on Textual Inversion, enabling privacy-preserving style transfer without the need for full model fine-tuning. By learning per-image embeddings and aggregating them with calibrated noise, our method, DPAgg-TI, achieves strong formal privacy guarantees while maintaining high output fidelity. Experiments on private artwork and Paris 2024 pictograms show that DPAgg-TI consistently outperforms DP-SGD, which fails to produce meaningful results under comparable privacy budgets. These results highlight the effectiveness of embedding-level adaptation as an efficient and scalable alternative to traditional gradient-based approaches, especially in low-data regimes. Unlike DP-SGD, which introduces significant computational overhead and utility degradation, DPAgg-TI is lightweight, modular, and compatible with existing diffusion backbones. Our findings suggest that embedding-centric approaches offer a promising direction for privacy-aware personalization, and motivate further research into cross-modal extensions, improved aggregation techniques, and integration with broader privacy-preserving frameworks.
Acknowledgements
We sincerely thank @eveismyname for providing her artwork for use in this study. We are also grateful to Anwar Hithnawi and Varun Chandrasekaran for their insightful discussions and feedback, as well as to all participants in our user study. Sanmi Koyejo acknowledges support by NSF 2046795 and 2205329, IES R305C240046, the MacArthur Foundation, Stanford HAI, OpenAI, and Google.
Ethical Statement
@@ Appendix A User Study
regular TI No Adaptation Unsure
@eveismyname 19 4 2
Paris 2024 16 6 3
DPAgg-TI (no DP, no subsampling) No Adaptation Unsure
@eveismyname 16 9 0
Paris 2024 15 4 6
regular TI DPAgg-TI (no DP, no subsamp.) Unsure
@eveismyname 12 13 0
Paris 2024 9 10 6
regular TI DPAgg-TI (no DP, subsamp. $m=8$) Unsure
@eveismyname 16 6 3
Paris 2024 7 13 5
@eveismyname 18 4 3
Paris 2024 10 8 7
@eveismyname 14 10 1
Paris 2024 3 16 6
DPAgg-TI (no DP, no subsampling) Style Guidance Unsure
@eveismyname 16 8 1
Paris 2024 20 2 3
DPAgg-TI ($\varepsilon=1$, subsamp. $m=8$) Style Guidance Unsure
@eveismyname 16 8 1
Paris 2024 19 2 4
@eveismyname 8 5 12
Paris 2024 15 4 6
Table 1: Survey Results.
@@ A.5 Results and Analysis
Appendix B Kernel Inception Distance
$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx 0$
0.0441 ± 0.0027 0.0798 ± 0.0032 0.0526 ± 0.0022 0.0688 ± 0.0020 0.1114 ± 0.0032 0.0654 ± 0.0027
32 0.0753 ± 0.0047 0.0836 ± 0.0042 0.1166 ± 0.0037 0.0295 ± 0.0019 0.0644 ± 0.0021 0.0650 ± 0.0025
16 0.0350 ± 0.0020 0.0381 ± 0.0018 0.0663 ± 0.0025 0.1303 ± 0.0033 0.0438 ± 0.0030 0.0660 ± 0.0029
8 0.0359 ± 0.0018 0.0364 ± 0.0017 0.0366 ± 0.0019 0.0394 ± 0.0025 0.0527 ± 0.0033 0.0654 ± 0.0024
4 0.0246 ± 0.0013 0.0251 ± 0.0016 0.0249 ± 0.0014 0.0256 ± 0.0012 0.0313 ± 0.0017 0.0653 ± 0.0023
ctrl 0.0314 ± 0.0010
Table 2: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl). Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx 0$
0.1153 ± 0.0055 0.1194 ± 0.0054 0.1306 ± 0.0046 0.1395 ± 0.0057 0.1201 ± 0.0053 0.1274 ± 0.0055
32 0.1222 ± 0.0066 0.1036 ± 0.0065 0.1375 ± 0.0047 0.1311 ± 0.0048 0.1248 ± 0.0060 0.1258 ± 0.0054
16 0.1321 ± 0.0057 0.1411 ± 0.0077 0.1309 ± 0.0061 0.1380 ± 0.0047 0.1359 ± 0.0060 0.1273 ± 0.0057
8 0.1303 ± 0.0084 0.1303 ± 0.0074 0.1112 ± 0.0062 0.1311 ± 0.0064 0.1318 ± 0.0052 0.1267 ± 0.0056
4 0.1158 ± 0.0057 0.1085 ± 0.0056 0.1184 ± 0.0068 0.1194 ± 0.0065 0.1592 ± 0.0065 0.1268 ± 0.0055
ctrl 0.1383 ± 0.0066
Table 3: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl). Reported values are the mean $\pm$ standard deviation over 100 random subsamples.
@@ Appendix B Kernel Inception Distance
For KID evaluations, we used prompts similar to those employed during TI training: “A painting/icon in the style of $S^{*}$”. Consistent with the training image captions, these prompts do not specify a subject. For each parameter configuration, we generate 100 images and compute KID by repeatedly subsampling the larger of the real and generated sets to match the size of the smaller set, 100 times, then averaging the resulting KID scores.
@@
Appendix C Differentially Private Adaptation via Style Guidance
@@ C.1 Background: Denoising Diffusion Implicit Models
Denoising Diffusion Implicit Models (DDIM) sampling (Song, Meng, and Ermon 2021) uses the predicted noise $\epsilon_{\theta}(x_{t},y,t)$ and a noise schedule represented by an array of scalars $\{\alpha_{t}\}_{t=1}^{T}$ to first predict a clean image $\hat{x}_{0}$, then makes a small step in the direction of $\hat{x}_{0}$ to obtain $x_{t-1}$. The reverse diffusion process for DDIM sampling can be formalized as follows: